# Proximal Oracles for Optimization and Sampling

### v2

@@ 4 Adaptive proximal bundle method
     if (19) is always true in the execution of Algorithm 2then
@@
Proof a) It follows from Lemma 3.1(d) that for every $u\in\mathbb{R}^{d}$
@@
d) This statement immediately follows from (b) and step 3 of Algorithm 3. ∎
@@
Assume $f$ is convex and $L_{\alpha}$-Hölder smooth. If $\eta_{0}\leq\|y_{0}-x_{*}\|^{2}/\varepsilon$, then the iteration-complexity to obtain an $\varepsilon$-solution to (1) is given by
@@
Proof Noting from Lemma 4.1 (d) that $\eta_{k}$ is bounded from below, we know there exists some $\tilde{\eta}\in[\underline{\eta},\eta_{0}]$ such that for some $k_{0}\geq 1$, $\eta_{k}\equiv\tilde{\eta}$ for $k\geq k_{0}$. Thus, it follows from the assumption that $\eta_{0}\leq\|y_{0}-x_{*}\|^{2}/\varepsilon$ that
@@
for cycles $k\leq k_{0}-1$. Finally, the total iteration-complexity (23) clearly follows from the above two bounds. ∎
@@ 5.1 Sampling from Hölder smooth potentials
  1. Run Algorithm 2 to compute $x_{J}$ and $\tilde{x}_{J}$.
@@
  if $U\leq\exp\left(-f_{y}^{\eta}(X)+h_{1}(X)\right)$then
@@
Proof The first inequality in (34) immediately follows from Lemma 3.1(d) and the definition of $h_{1}$ in (33a). By the definition of $f_{y}^{\eta}$ in (9) we get
@@
Using the above inequality and the definition of $h_{2}$ in (33b), we conclude that the second inequality in (34) holds. ∎
@@
Proof Denote $r=\|x\|$, then
@@
where $\mathrm{d}S^{d-1}$ is the surface area of the $(d-1)$-dimensional unit sphere. It follows that
@@
In the above equation, we have used the fact that the total surface area of a $(d-1)$-dimensional unit sphere is $2\pi^{d/2}/\Gamma\left(\frac{d}{2}\right)$ where $\Gamma(\cdot)$ is the gamma function, i.e.,
@@
The result (37) then follows from the above inequality and (38). ∎
@@
Proof It is a well-known result for rejection sampling that $X\sim\pi^{X|Y}(x\mid y)$ and the probability that $X$ is accepted is
@@
Proof The results follow directly from Theorem 5.1, Theorem 3.5 and Proposition 5.4 with the choice of stepsize $\eta\asymp 1/(L_{\alpha}^{\frac{2}{\alpha+1}}d)$. ∎
@@ 5.2 Sampling from hybrid potentials
Proof Using the Young’s inequality $st\leq s^{p}/p+t^{q}/q$ with
@@
where the second inequality is due to the fact that for $\alpha\in[0,1]$, $(\alpha+1)2^{\frac{-2\alpha}{\alpha+1}}\leq 1$. Hence, the above inequality generalizes to
@@ 6 Conclusions
This paper provides a unified perspective to study proximal optimization and sampling algorithms, while many other interesting questions remain open. First, APBM is only optimal when $\alpha=0$, i.e., $f$ is Lipschitz continuous. We are interested in developing a universal method that is optimal for any $\alpha\in[0,1]$. One possible direction is to incorporate the acceleration technique into both the regularized cutting-plane method and the PPF. Second, as acceleration methods are widely used in optimization to obtain optimal performance, accelerated proximal sampling algorithms are less explored. It is worth investigating a counterpart of the accelerated proximal point method [42] in sampling. Finally, we develop APBM as a universal method for non-smooth optimization, it would be equally important to design a universal method for sampling.
Acknowledgement
This work was supported by NSF under grants 1942523 and 2008513.
@@ Appendix A Technical results
For $10<s<$ 1 and $t>0$, the gamma function defined as in (39) satisfies

### v3

@@ 4 Adaptive Proximal Bundle Method
  if (21) is always true in the execution of Algorithm 2, then
@@
Proof: a) It follows from Lemma 3.1(d) that for every $u\in\mathbb{R}^{d}$
@@
d) This statement immediately follows from (b) and step 3 of Algorithm 3.
@@
Assume $f$ is convex and $L_{\alpha}$-Hölder smooth. If $\eta_{0}\leq\|y_{0}-x_{*}\|^{2}/\varepsilon$, then the iteration-complexity to obtain an $\varepsilon$-solution to (1) (i.e., a point $\hat{x}$ such that $f(\hat{x})-\min_{x\in\mathbb{R}^{d}}f(x)\leq\varepsilon$) is given by
@@
Proof: Noting from Lemma 4.1 (d) that $\eta_{k}$ is bounded from below, we know there exists some $\tilde{\eta}\in[\underline{\eta},\eta_{0}]$ such that for some $k_{0}\geq 1$, $\eta_{k}\equiv\tilde{\eta}$ for $k\geq k_{0}$. Thus, it follows from the assumption that $\eta_{0}\leq\|y_{0}-x_{*}\|^{2}/\varepsilon$ that
@@
for cycles $k\leq k_{0}-1$. Finally, the total iteration-complexity (25) clearly follows from the above two bounds.
We note that the final $\varepsilon$-solution produced by Algorithm 3 is the point $\tilde{y}_{k}$ that achieves $\min_{1\leq k\leq n}f(\tilde{y}_{k})$. This differs from the last-iterate convergence observed in the smooth case, since the objective function $f$ here is Hölder smooth and may include the nonsmooth case (i.e., $\alpha=0$).
@@ 5.1 Sampling from Hölder Smooth Potentials
 1. Let $y\in\mathbb{R}^{d}$, $\eta>0$, and $\delta>0$ be given, and run Algorithm 2 to compute $x_{J}$ and $\tilde{x}_{J}$.
@@
if $U\leq\exp\left(-f_{y}^{\eta}(X)+h_{1}(X)\right)$, then
@@
Proof: The first inequality in (36) immediately follows from Lemma 3.1(d) and the definition of $h_{1}$ in (35a). By the definition of $f_{y}^{\eta}$ in (11) we get
@@
Using the above inequality and the definition of $h_{2}$ in (35b), we conclude that the second inequality in (36) holds.
@@
Proof: Denote $r=\|x\|$, then
@@
where $\mathrm{d}S^{d-1}$ is the surface area of the (d1)(d-1)-dimensional unit sphere. It follows that
@@
In the above equation, we have used the fact that the total surface area of a (d1)(d-1)-dimensional unit sphere is $2\pi^{d/2}/\Gamma\left(\frac{d}{2}\right)$ where $\Gamma(\cdot)$ is the gamma function, i.e.,
@@
The result (39) then follows from the above inequality and (40).
@@
Proof: It is a well-known result for rejection sampling that $X\sim\pi^{X|Y}(x\mid y)$ and the probability that $X$ is accepted is
@@
Proof: The results follow directly from Theorem 5.1, Theorem 3.5 and Proposition 5.4 with the choice of stepsize $\eta\asymp 1/(L_{\alpha}^{\frac{2}{\alpha+1}}d)$.
@@ 5.2 Sampling from Hybrid Potentials
Proof: Using the Young’s inequality $st\leq s^{p}/p+t^{q}/q$ with
@@
where the second inequality is due to the fact that $(\alpha+1)2^{\frac{-2\alpha}{\alpha+1}}\leq 1$ for $\alpha\in[0,1]$. Hence, the above inequality generalizes to
@@ 6 Conclusions
This paper provides a unified perspective to study proximal optimization and sampling algorithms, while many other interesting questions remain open. First, APBM is only optimal when $\alpha=0$, i.e., $f$ is Lipschitz continuous. We are interested in developing a universal method that is optimal for any $\alpha\in[0,1]$. One possible direction is to incorporate the acceleration technique into both the regularized cutting-plane method and the PPF. Second, as acceleration methods are widely used in optimization to obtain optimal performance, accelerated proximal sampling algorithms are less explored. It is worth investigating a counterpart of the accelerated proximal point method [46] in sampling. Finally, we develop APBM as a universal method for non-smooth optimization, and it would be equally important to design a universal method for sampling.
@@ Appendix A Technical results
For $0<s<1$ and $t>0$, the gamma function defined as in (41) satisfies
@@
Lemma A.4.
Consider $\phi(t)=|t|^{p}$ and $\phi^{\prime}(t)=p\,\operatorname{sign}(t)\,|t|^{p-1}$ for $t\in\mathbb{R}$ and some $p\in[1,2]$. Then, for any $u,v\in\mathbb{R}$, we have
$|\phi^{\prime}(u)-\phi^{\prime}(v)|\leq p\,2^{2-p}\,|u-v|^{p-1}.$
Proof: Let $r:=p-1\in[0,1]$. We consider the following two cases and prove
$|\phi^{\prime}(u)-\phi^{\prime}(v)|\leq p\,2^{1-r}\,|u-v|^{r}.$
Case 1: $uv\geq 0$. Here, $\operatorname{sign}(u)=\operatorname{sign}(v)$, so
$|\phi^{\prime}(u)-\phi^{\prime}(v)|=p\,\big||u|^{r}-|v|^{r}\big|.$
Without loss of generality, assume $a=|u|\geq b=|v|\geq 0$. By the subadditivity of the function $x\mapsto x^{r}$ with $r\in[0,1]$, we have $(a-b)^{r}\geq a^{r}-b^{r}$ for all $a\geq b\geq 0$. Hence
$|\,|u|^{r}-|v|^{r}\,|=a^{r}-b^{r}\leq(a-b)^{r}=\big||u|-|v|\big|^{r}=|u-v|^{r}.$
Therefore,
$|\phi^{\prime}(u)-\phi^{\prime}(v)|\leq p\,|u-v|^{r}.$
Case 2: $uv<0$. In this case, $\operatorname{sign}(u)=-\operatorname{sign}(v)$, so
$|\phi^{\prime}(u)-\phi^{\prime}(v)|=p\,(|u|^{r}+|v|^{r}).$
By the concavity of $x\mapsto x^{r}$, we have
$|u|^{r}+|v|^{r}\leq 2^{1-r}(|u|+|v|)^{r}=2^{1-r}|u-v|^{r},$
and thus
$|\phi^{\prime}(u)-\phi^{\prime}(v)|\leq p\,2^{1-r}\,|u-v|^{r}.$
The conclusion immediately follows from the above two cases.